a man had two sons, albert and bernie, their combined age was 30 and albert is 4 years older than bernie
Answer by macston(5194)

Albert's age =A, A=B+4; Bernie's age=B
A+B=30 Substitute for A
B+4+B=30
2B+4=30 Subtract 4 from each side
2B=26 Divide each side by 2
B=13 ANSWER 1: Bernie is 13 years old
A=B+4=13+4=17 ANSWER 2: Albert is 17 years old
CHECK
A+B=30
13+17=30
30=30